OpenStudy (anonymous):

find the slope (2,-3) and (4,-1)

OpenStudy (a_clan):

2/2 =1

OpenStudy (anonymous):

-7/3

OpenStudy (anonymous):

how'd you get -7/3?

OpenStudy (anonymous):

there is a formula:\[y _{2}-y _{1}/x _{2}-x _{1}\]

OpenStudy (radar):

slope (m)= (y2-y1)/(x2-x1) Let us call (2,-3) P1 and Let us call (4, -1) P2. Please note that this is your call and it is arbitrary but once called, you must stick with it. m=(-1--3)/(4-2) m=(2/2)=1 m=1

OpenStudy (radar):

Did you follow with understanding?